# ViCTr: Vital Consistency Transfer for Pathology Aware Image Synthesis

### v2

@@ 1 Introduction
Figure 1: Overview of the proposed ViCTr methodology. Snowflake symbol () indicates frozen weights, and the fire symbol denotes trainable weights.
@@ 2.2 Diffusion Models in Medical Imaging
Diffusion-based approaches have emerged as a powerful alternative to GANs for high-quality synthetic image generation [49, 4, 5, 13]. Their inherent noise-to-image paradigm provides a more stable training regime and can produce richer data variations. In medical imaging, diffusion models have been leveraged to improve segmentation and classification performance by generating diverse training samples [46, 10, 47, 37, 1]. Recent specialized frameworks, including segmentation-guided diffusion [26] and text-driven generation built upon RadImageNet [54, 34], highlight the adaptability of diffusion models to specific clinical scenarios. Efforts such as MixUp-enhanced augmentation [28, 6] and domain adaptation methods [55] demonstrate that diffusion techniques can mitigate biases and improve generalization in tasks like classification and risk prediction.
@@ 2.5 Domain-Specific Adaptation
Figure 2: Comparison of synthetic data distributions generated by various methods against the ground-truth (GT) distribution on the AMOS dataset. The GT distribution (blue) represents the real data distribution, while the synthetic distributions generated by Reflow (red), Flow Straight and Fast (green), and Distribution Matching Distillation (purple) demonstrate varying degrees of alignment. Our proposed method, ViCTr (brown), achieves the closest match to the GT distribution, showcasing superior distribution alignment and data fidelity.
@@ 3 Methods
Our proposed framework, ViCTr, leverages Rectified Flow [31] and Tweedie’s Formula [43] to synthesize high-fidelity medical images with efficient, one-step sampling. By rectifying the trajectory from a prior distribution $p_{0}$ to the target medical image distribution $p_{\text{target}}$, we reduce sampling bias and preserve crucial anatomical and pathological details.
@@ 3.1 Rectified Flow Trajectory
encourages $\hat{\mathcal{T}(x_{0})}$ to approximate the final, rectified state $x_{1}$.
@@ 3.2 Rectified Flow with Tweedie’s Formula
While rectified flow refines the sampling path, medical data distributions often require additional bias correction. Tweedie’s Formula [43] addresses this by adjusting noisy observations to better approximate the posterior mean. For Gaussian variables $z\sim\mathcal{N}(\mu_{z},\Sigma_{z})$, Tweedie’s formula indicates:
@@
, where $\nabla_{z}\log p(z)$ is the gradient of the log probability. In typical diffusion, Tweedie’s formula defines the estimate of $x_{0}$ from noisy observation $x_{t}$, guiding the predicted states closer to the data manifold.
@@ 3.3 Pathology Aware Image Synthesis
In the absence of large-scale, domain-specific pre-trained diffusion models, we begin by training on the ATLAS-8k [40] dataset, which comprises abdominal CT scans and their segmentation annotations. We leverage these annotations as conditional guidance—enabling precise control over anatomical structures—and integrate textual prompts (e.g., “create image having <<much-less-than$<$ <organs>>much-greater-than$>$ >) to further refine semantic consistency.
@@
Building on the pre-trained model from Stage 1, Stage 2 focuses on specialized applications such as semantic-guided CT/MRI generation and counterfactual pathology synthesis. As shown in Figure 1, we employ a dual-network arrangement:
@@
Training Input and Consistency Loss. Similar to Stage 1, each training sample includes prompts, segmentation maps, and either CT or MRI images (depending on the task). A frozen VAE encoder extracts latent representations $Z_{s},Z_{o}$, and $Z_{p}$. Following forward diffusion of $Z_{o}$ into noisy $Z_{t}$, both $\phi_{base}$ and $\phi_{adapt}$ process $[Z_{t},Z_{s}]$ with text embedding $Z_{p}$. We introduce a consistency loss $L_{consistency}$ that aligns intermediate outputs of $\phi_{base}$ and $\phi_{adapt}$, promoting stable adaptation without deviating from foundational representations.
@@
Pathology Generation. For liver cirrhosis, we employ CirrMRI600+ dataset (T1/T2 MRIs) with segmentation masks marking liver regions. Prompts specify “low,” “mild,” or “severe” cirrhosis intensity. As in the CT/MRI setting, the training/validation/testing split is 80/10/10. The inference details and inference diagram are in the supplementary material.
@@ 4.1 Quantitative Results
Baselines BTCV Dataset (CT Generation) AMOS Dataset (MRI Generation) CirrMRI600+ (Pathology Generation) Diffusion Steps |||| Inference Time
@@
Stable Diffusion [44] 25.44 / 19.67 21.98 / 19.02 25.43 / 21.76 20.37 / 19.11 28.34 / 23.43 25.57 / 21.46 40 |||| 13.76 4 |||| 3.12
Stable Diffusion XL [39] 23.47 / 18.21 20.33 / 17.44 24.11 / 20.23 19.44 / 18.45 27.34 / 22.11 24.02 / 20.76 30 |||| 14.55 4 |||| 3.45
Stable Diffusion-3 [14] 19.07 / 16.22 17.37 / 16.02 22.32 / 19.76 18.02 / 19.08 24.49 / 21.78 21.28 / 19.34 50 |||| 18.98 3 |||| 2.78
Pixart-alpha [9] 21.32 / 17.09 19.22 / 16.96 23.78 / 20.04 18.76 / 18.56 26.06 / 20.07 23.04 / 18.92 25 |||| 10.67 3 |||| 1.74
Flux [27] 15.52 / 15.01 13.28 / 14.08 19.02 / 18.28 15.55 / 16.58 22.46 / 18.88 19.96 / 17.01 30 |||| 15.66 3 |||| 2.87
Table 2: Quantitative results on CT, MRI, and Cirrhosis generation. All baselines were pre-trained on ATLAS-8k and fine-tuned on each target dataset. The reported metrics are Fréchet Inception Distance (FID) and Medical Fréchet Inception Distance (MFID)-separated by ”/”, and inference time is measured in seconds. Lower values of these metrics indicate better performance.
@@
Table 3: Segmentation results on BTCV, AMOS, and CirrMRI600+ datasets with different training settings and baseline models. Metrics reported are mDSC (%) and mHD95 (mm). The top three results for each setting are highlighted: blue for 1st, green for 2nd, and orange for 3rd.
@@
Segmentation Results. To assess the quality and utility of ViCTr-generated synthetic data, we performed segmentation experiments on three datasets—BTCV, AMOS, and CirrMRI600+—tracking mean Dice Similarity Coefficient (mDSC) and mean Hausdorff Distance 95 (mHD95). Higher mDSC and lower mHD95 respectively indicate improved overlap accuracy and spatial precision. We examined four training configurations: (1) baseline using original data only, (2) standard augmentation (Random Crop, Rotate, Blur, Affine, Geometric Distortion at 0.4 probability), (3) 30% synthetic data via vanilla fine-tuning, and (4) 30% synthetic data from ViCTr. Table 3 highlights consistent performance gains from ViCTr across all datasets and backbone architectures. Notably, CirrMRI600+—a liver cirrhosis dataset—exhibited the most significant improvements, reflecting ViCTr’s capacity to capture complex pathological features. These results confirm that ViCTr-generated synthetic images can substantially enhance segmentation metrics (mDSC and mHD95) in medical imaging tasks, underscoring the practical value of our two-stage framework for data augmentation.
@@ 4.2 Qualitative Results
We conducted extensive qualitative assessments of ViCTr across multiple imaging modalities (CT and MRI) and tasks (including both anatomical and pathological image synthesis).
Anatomical and Pathology-Driven Image Generation. Unlike prior methods that primarily generate normal anatomical scans or isolated tumors, ViCTr extends synthesis capabilities to complex pathologies, such as liver cirrhosis. Figure 5 shows pairs of non-cirrhotic MRIs alongside their segmentation masks and the corresponding ViCTr-generated cirrhotic images. Notably, the synthesized images retain essential anatomical structures while introducing realistic cirrhotic texturing, an advancement that enables more diverse data augmentation and supports research into disease progression. These high-fidelity synthetic samples hold promise for enhancing clinical training datasets, improving diagnostic algorithm development, and facilitating in-depth analyses of pathology evolution.
Qualitative Segmentation Outcomes. To further demonstrate ViCTr’s impact on downstream tasks, we visually evaluated segmentation quality on the BTCV and AMOS datasets (Figure6). While these datasets encompass multiple abdominal organs, our illustrations highlight selected regions such as the left kidney and pancreas. Models trained with ViCTr-augmented data exhibit sharper organ boundaries and more precise delineations compared to standard fine-tuning methods, indicating enhanced spatial learning.
@@
Second, we evaluate the impact of ViCTr on segmentation tasks, illustrated in Figure 6. Both BTCV and AMOS datasets encompass various abdominal structures (e.g., left kidney, pancreas), and our results highlight how ViCTr-based training yields notably more precise organ boundaries compared to alternative methods. This improvement is particularly evident in challenging regions where standard segmentation approaches often struggle, further demonstrating ViCTr’s effectiveness at enhancing spatial delineation and anatomical accuracy.
@@ 4.3 Ablation Studies
Ablation Study Based on Loss Functions. We then investigated each ViCTr loss component—diffusion loss, spatial consistency, and consistency loss—to determine their relative contributions: the baseline model, utilizing only diffusion loss, produced an FID of $18.77$ (Table 4). Incorporating spatial consistency loss yielded a moderate improvement to $18.21$, highlighting its role in maintaining structural integrity. The addition of consistency loss further enhanced performance, reducing the FID to $17.02$ by promoting image coherence across generations. When combining all three terms, we achieved an FID of 15.55, indicating that these components collectively offer significant improvements in anatomical accuracy and visual realism.
Ablation Based on LoRA Rank. Finally, we explored how varying the LoRA rank ($r$) influences generative performance. Starting at $r=8$ and incrementing upwards, we found $r=64$ to yield the best FID at $15.55$. This result suggests that higher-dimensional adaptation spaces enable more nuanced parameter updates during fine-tuning, thereby improving visual quality and fidelity.
Visual Turing Tests. To further validate clinical plausibility, three radiologists participated in Visual Turing Tests using 15 randomly generated MRI scans depicting varying levels of liver cirrhosis (mild, moderate, severe). All scans were uniformly judged to be clinically realistic, with identical outcomes observed when using additional random samples. Notably, the synthetic cirrhotic images correctly exhibited surface nodularity and textural irregularities, consistent with radiologic findings in mild-to-severe cirrhosis (Figure 5). These results confirm ViCTr’s capacity to synthesize pathology-specific features, supporting potential applications in training, algorithm development, and clinical research.
@@ 5 Conclusion
We presented ViCTr, a novel two-stage framework designed to generate high-fidelity medical images by integrating robust anatomical pre-training with precise pathology-specific fine-tuning. By aligning Tweedie’s formula with linear projection methods used in flow matching, ViCTr maintains accurate initial distribution estimates even amid diffusion processes. This setup is further enhanced by LoRA adapters, which preserve essential anatomical information while flexibly adapting to diverse pathologies. Experimental results demonstrate that ViCTr consistently surpasses conventional fine-tuning strategies, reducing FID scores and producing clinically realistic outputs. Beyond data augmentation for segmentation and classification, ViCTr can be extended to tasks such as modality translation, contrast synthesis, and professional training, offering a powerful and versatile tool for advancing AI-driven medical imaging.  [^1] We will publicly release all models, code, and preprocessed datasets. We are working with HF-Community to release it with HF-Diffusers
@@ 6 Quantitative
Additional Segmentation Results: We present extended visual results showcasing segmentation performance on complex organs such as the spleen, liver, aorta, and stomach. As depicted in Figure 11, our method, which leverages synthetic data generated via the Flux (ViCtr) framework, demonstrates superior alignment with ground truth (GT) segmentation. Notably, the quality and consistency of the predicted masks across all four organ classes are on par with GT annotations. These results highlight the efficacy of our approach in capturing intricate organ structures with high precision and robustness.
7 Qualitative
Figure 9: Synthetically generated MRI images using ViCTr.
8 Training and Implementation Details
Pre-training. We pre-trained ViCTr Stage-1 using a rectified flow strategy, with the maximum diffusion steps set to 100. The Atlas-8K dataset was used as the foundational dataset, and training was performed at an image resolution of $256\times 256$. We employed a batch size of $8$, with gradient accumulation over 8 steps. Optimization was carried out using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, managed by a cosine annealing scheduler to ensure a smooth decay of the learning rate over time. The pre-training phase was conducted on 4 nodes, each equipped with 8 Nvidia A100 GPUs 80GB each, and completed in approximately 52 hours.

### v3

@@ 1 Introduction
Figure 1: Overview of the proposed ViCTr methodology.
@@ 2.2 Diffusion Models in Medical Imaging
Diffusion-based approaches have emerged as a powerful alternative to GANs for high-quality synthetic image generation [49, 4, 5, 13]. Their inherent noise-to-image paradigm provides a more stable training regime and can produce richer data variations. In medical imaging, diffusion models [11] have been leveraged to improve segmentation and classification performance by generating diverse training samples [46, 10, 47, 37, 1]. Recent specialized frameworks, including segmentation-guided diffusion [26] and text-driven generation built upon RadImageNet [54, 34], highlight the adaptability of diffusion models to specific clinical scenarios. Efforts such as MixUp-enhanced augmentation [28, 6] and domain adaptation methods [55] demonstrate that diffusion techniques can mitigate biases and improve generalization in tasks like classification and risk prediction.
@@ 2.5 Domain-Specific Adaptation
Figure 2: Comparison of synthetic data distributions from various methods against the ground-truth (GT) on the AMOS dataset. The GT distribution (blue) reflects real data, while Reflow (red), Flow Straight and Fast (green), and Distribution Matching Distillation (purple) show varying alignment. Our proposed method, ViCTr (brown) achieves the closest match to GT, demonstrating superior distribution alignment and fidelity.
@@ 3 Methods
Our framework, ViCTr, uses Rectified Flow [31] and Tweedie’s Formula [43] for high-fidelity medical image synthesis with efficient one-step sampling. By rectifying the trajectory from a prior distribution $p_{0}$ to the target medical image distribution $p_{\text{target}}$, we reduce sampling bias and retain key anatomical and pathological details.
@@ 3.1 Rectified Flow Trajectory
drives $\hat{\mathcal{T}(x_{0})}$ to approximate the final, rectified state $x_{1}$.
@@ 3.2 Rectified Flow with Tweedie’s Formula
Though rectified flow refines sampling, medical data distributions often need extra bias correction. Tweedie’s Formula [43] addresses this by adjusting noisy observations to better approximate the posterior mean. For Gaussian variables $z\sim\mathcal{N}(\mu_{z},\Sigma_{z})$, Tweedie’s formula indicates:
@@
, where $\nabla_{z}\log p(z)$ is the gradient of the log probability. In typical diffusion, Tweedie’s formula estimates $x_{0}$ from noisy $x_{t}$, guiding predictions toward the data manifold.
@@ 3.3 Pathology Aware Image Synthesis
In the absence of large-scale, domain-specific pre-trained diffusion models, we begin by training on the ATLAS-8k [40] dataset, which comprises abdominal CT scans and their segmentation annotations. We leverage these annotations as conditional guidance—enabling precise control over anatomical structures—and integrate textual prompts (e.g., “create image having <<<<< <organs>>>>> >”) to further refine semantic consistency.
@@
Building on the Stage1 pre-trained model, Stage2 targets specialized tasks like semantic-guided CT/MRI generation and counterfactual pathology synthesis. As shown in Figure 1, a dual-network setup is adopted.
@@
Training Input & Consistency Loss. Similar to Stage 1, each training sample includes prompts, segmentation maps, and either CT or MRI images (depending on the task). A frozen VAE encoder extracts latent representations $Z_{s},Z_{o}$, and $Z_{p}$. Following forward diffusion of $Z_{o}$ into noisy $Z_{t}$, both $\phi_{base}$ and $\phi_{adapt}$ process $[Z_{t},Z_{s}]$ with text embedding $Z_{p}$. We introduce a consistency loss $L_{consistency}$ to align intermediate outputs of $\phi_{base}$ and $\phi_{adapt}$, ensuring stable adaptation without drifting from core representations.
@@
Pathology Generation. For liver cirrhosis, we employ CirrMRI600+ [22] dataset (T1/T2 MRIs) with segmentation masks marking liver regions. Prompts specify “low,” “mild,” or “severe” cirrhosis intensity. As in the CT/MRI setting, the training/validation/testing split is 80/10/10. The inference details and inference diagram are in the supplementary material.
@@ 4.1 Quantitative Results
Baselines BTCV Dataset (CT Generation) AMOS Dataset (MRI Generation) CirrMRI600+ (Pathology Generation) Diffusion Steps ||| Inference Time
@@
Stable Diffusion [44] 25.44 / 19.67 21.98 / 19.02 25.43 / 21.76 20.37 / 19.11 28.34 / 23.43 25.57 / 21.46 40 ||| 13.76 4 ||| 3.12
Stable Diffusion XL [39] 23.47 / 18.21 20.33 / 17.44 24.11 / 20.23 19.44 / 18.45 27.34 / 22.11 24.02 / 20.76 30 ||| 14.55 4 ||| 3.45
Stable Diffusion-3 [14] 19.07 / 16.22 17.37 / 16.02 22.32 / 19.76 18.02 / 19.08 24.49 / 21.78 21.28 / 19.34 50 ||| 18.98 3 ||| 2.78
Pixart-alpha [9] 21.32 / 17.09 19.22 / 16.96 23.78 / 20.04 18.76 / 18.56 26.06 / 20.07 23.04 / 18.92 25 ||| 10.67 3 ||| 1.74
Flux [27] 15.52 / 15.01 13.28 / 14.08 19.02 / 18.28 15.55 / 16.58 22.46 / 18.88 19.96 / 17.01 30 ||| 15.66 3 ||| 2.87
Table 2: Quantitative results on CT, MRI, and Cirrhosis generation. All baselines were pre-trained on ATLAS-8k and fine-tuned on target datasets. Metrics reported are Fréchet Inception Distance (FID) and Medical FID (MFID), shown as FID/MFID, with inference time in seconds. Lower values indicate better performance.
@@
Table 3: Segmentation results on BTCV, AMOS, and CirrMRI600+ using various training settings and baselines. Metrics reported are mDSC (%) and mHD95 (mm). Top three results per setting are highlighted: blue (1st), green (2nd), and orange (3rd).
@@
Segmentation Results. To assess the quality and utility of ViCTr-generated synthetic data, we performed segmentation experiments on three datasets—BTCV, AMOS, and CirrMRI600+—tracking mean Dice Similarity Coefficient (mDSC) and mean Hausdorff Distance 95 (mHD95). Higher mDSC and lower mHD95 respectively indicate improved overlap accuracy and spatial precision. We examined four training configurations: (1) baseline using original data only, (2) standard augmentation (Random Crop, Rotate, Blur, Affine, Geometric Distortion at 0.4 probability), (3) 30% synthetic data via vanilla fine-tuning, and (4) 30% synthetic data from ViCTr. Table 3 highlights consistent performance gains from ViCTr across all datasets and backbone architectures. Notably, CirrMRI600+—a liver cirrhosis dataset—exhibited the most significant improvements, reflecting ViCTr’s capacity to capture complex pathological features. These results confirm that ViCTr-generated images significantly boost segmentation metrics (mDSC and mHD95), highlighting the practical value of our two-stage framework for medical data augmentation.
@@ 4.2 Qualitative Results
We conducted extensive qualitative evaluations of ViCTr across multiple modalities (CT and MRI) and tasks, covering both anatomical and pathological image synthesis.
Anatomical and Pathology-Driven Image Generation. Unlike prior methods that primarily generate normal anatomical scans or isolated tumors, ViCTr extends synthesis capabilities to complex pathologies, such as liver cirrhosis. Figure 5 shows pairs of non-cirrhotic MRIs alongside their segmentation masks and the corresponding ViCTr-generated cirrhotic images. Notably, the synthesized images retain essential anatomical structures while introducing realistic cirrhotic texturing, an advancement that enables more diverse data augmentation and supports research into disease progression. These high-fidelity synthetic samples offer potential to enrich clinical training datasets, advance diagnostic algorithms, and support detailed analyses of pathology progression.
Qualitative Segmentation Outcomes. To further showcase ViCTr’s impact on downstream tasks, we visually assessed segmentation quality on BTCV and AMOS (Figure 6). Focusing on regions like the left kidney and pancreas, models trained with ViCTr-augmented data show sharper boundaries and more precise delineation than standard fine-tuning, indicating improved spatial learning.
@@
Second, we evaluate the impact of ViCTr on segmentation tasks, illustrated in Figure 6. Both BTCV and AMOS datasets encompass various abdominal structures (e.g., left kidney, pancreas), and our results highlight how ViCTr-based training yields notably more precise organ boundaries compared to alternative methods. This improvement is particularly evident in challenging regions where standard segmentation often struggle, highlighting ViCTr’s strength in enhancing spatial delineation and anatomical accuracy.
@@ 4.3 Ablation Studies
Ablation Study Based on Loss Functions. We then investigated each ViCTr loss component—diffusion loss, spatial consistency, and consistency loss—to determine their relative contributions: the baseline model, utilizing only diffusion loss, produced an FID of 18.7718.7718.77 (Table 4). Incorporating spatial consistency loss yielded a moderate improvement to 18.2118.2118.21, highlighting its role in maintaining structural integrity. The addition of consistency loss further enhanced performance, reducing the FID to 17.0217.0217.02 by promoting image coherence across generations. When combining all three terms, we achieved an FID of 15.55, indicating that these components collectively offer significant improvements in anatomical accuracy and visual realism.
Ablation Based on LoRA Rank. Finally, we explored how varying the LoRA rank ($r$) influences generative performance. Starting at $r=8$ and incrementing upwards, we found $r=64$ to yield the best FID at 15.5515.5515.55. This result suggests that higher-dimensional adaptation spaces enable more nuanced parameter updates during fine-tuning, thereby improving visual quality and fidelity.
Visual Turing Tests. To further validate clinical plausibility, three radiologists participated in Visual Turing Tests using 15 randomly generated MRI scans depicting varying levels of liver cirrhosis (mild, moderate, severe). All scans were uniformly judged to be clinically realistic, with identical outcomes observed when using additional random samples. Notably, the synthetic cirrhotic images correctly exhibited surface nodularity and textural irregularities, consistent with radiologic findings in mild-to-severe cirrhosis (Figure 5). These results confirm ViCTr’s ability to synthesize pathology-specific features, enabling applications in training, algorithm development, and clinical research.  [^1] Code: https://github.com/Onkarsus13/ViCTr-2D
Weights: https://huggingface.co/onkarsus13/ViCTr-2D
@@ 5 Conclusion
We presented ViCTr, a novel two-stage framework designed to generate high-fidelity medical images by integrating robust anatomical pre-training with precise pathology-specific fine-tuning. By aligning Tweedie’s formula with linear projection methods used in flow matching, ViCTr maintains accurate initial distribution estimates even amid diffusion processes. This setup is further enhanced by LoRA adapters, which preserve essential anatomical information while flexibly adapting to diverse pathologies. Experimental results demonstrate that ViCTr consistently surpasses conventional fine-tuning strategies, reducing FID scores and producing clinically realistic outputs. Beyond data augmentation for segmentation and classification, ViCTr can be extended to tasks such as modality translation, contrast synthesis, and professional training, offering a powerful and versatile tool for advancing AI-driven medical imaging.
Acknowledgments
This research is supported by the following NIH grants: R01-HL171376 and U01-CA268808.
@@ A Quantitative
Additional Segmentation Results:
Model Vanilla FineTuning ViCTr (Ours)
Stable Diffusion 23.11 / 84.72 25.27 / 86.72
Stable Diffusion-XL 24.34 / 85.72 26.78 / 87.93
Stable Diffusion-3 26.44 / 87.51 28.92 / 90.37
Pixart 26.32 / 88.21 31.09 / 91.33
Flux 27.51 / 90.21 33.33 / 94.05
Table 5: PSNR / SSIM (%) comparison between Vanilla FineTuning and our ViCTr across diffusion models on CirrMRI600+.
We present extended visual results showcasing segmentation performance on complex organs such as the spleen, liver, aorta, and stomach. As depicted in Figure 11, our method, which leverages synthetic data generated via the Flux (ViCtr) framework, demonstrates superior alignment with ground truth (GT) segmentation. Notably, the quality and consistency of the predicted masks across all four organ classes are on par with GT annotations. These results highlight the efficacy of our approach in capturing intricate organ structures with high precision and robustness.
Modality Translation Results on CirrMRI600+
Experimental Setup
To evaluate cross-modality translation performance, see in Table 5, we conducted experiments using the paired T1–T2 volumes from the CirrMRI600+ dataset. The goal was to synthesize target modality (T2-weighted) images conditioned on anatomical features from the source modality (T1-weighted) using text-based prompts such as “Generate the pathology on T2-weighted MRI”.
We assessed both structural preservation and pathological fidelity of the translated outputs. Quantitative evaluation was carried out using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM), comparing synthesized T2 volumes against ground truth.
Results
Our proposed ViCTr framework consistently outperformed baseline diffusion-based models across all metrics, demonstrating superior anatomical consistency and modality-specific detail reconstruction. These findings emphasize the potential of ViCTr in downstream clinical applications such as modality harmonization, synthetic augmentation, and diagnostic support.